Find the distance between the pair of parallel lines, y = -2x+1 & y = -2x+16.
Answer by lwsshak3(11628)

Find the distance between the pair of parallel lines, y = -2x+1 & y = -2x+16. (slope,m=-2)
equation of a line perpendicular to pair of given lines: y=x/2 (slope,m=1/2)
Point of intersection between perpendicular line and one of the parallel lines:
x/2=-2x+1
x=-4x+2
5x=2
x=2/5
y=1/5
point of intersection: (2/5,1/5)
..
Point of intersection between perpendicular line and other parallel line:
x/2=-2x+16
x=-4x+32
5x=32
x=32/5
y=32/10=16/5
point of intersection: (32/5,16/5)
..
Use distance formula to find the distance between the 2 points of intersection
d=√(x1-x2)^+(y1-y2)^2)=√(32/5-2/5)^2+(16/5-(1/5)^2=√(36+9)=√45
distance between the pair of parallel lines=√45